OpenStudy (anonymous):

What is the equation of a line passing through (-6, 5) and having a slope of 1/3?

OpenStudy (anonymous):

y=1/3x+5

OpenStudy (anonymous):

that's not one of my listed answers would it be y=1/3x+7

OpenStudy (anonymous):

First of all, remember what the equation of a line is: y = mx+b Where: m is the slope, and b is the y-intercept

OpenStudy (anonymous):

let me see all the listed answers

OpenStudy (anonymous):

The equation of the line that passes through the point (-6,5) with a slope of 1/3 is y=1/3x+7

OpenStudy (anonymous):

so far so good @pinkbubbles

OpenStudy (anonymous):

but where did the seven come from?

OpenStudy (anonymous):

(-6,5). y=mx+b or 5=1/3 × -6+b, or solving for b: b=5-(1/3)(-6). b=7

OpenStudy (anonymous):

@DaBest21

OpenStudy (anonymous):

A.y=1/3x-23/3 B.y=1/3x-3 C.y=1/3x+7 D.y=3x+7

OpenStudy (anonymous):

The equation of the line that passes through the point (-6,5) with a slope of 1/3 is y=1/3x+7

OpenStudy (anonymous):

ok thank u :D

OpenStudy (anonymous):

Welcome! @sherlocked1270 :)